OpenStudy (anonymous):

x+(2x+0.15) + (x+0.05)=1.8

OpenStudy (ahsome):

Equation: \[x+(2x+0.15) + (x+0.05)=1.8\]Right?

OpenStudy (ahsome):

Hello, @jennydoe? Is that the equation?

OpenStudy (ahsome):

Assuming it is:\[x+(2x+0.15) + (x+0.05)=1.8\] Brackets do not impact in this equation:\[x+2x+0.15+x+0.05)=1.8\] Simplify\[4x+0.2=1.8\] Move the \(0.2\) to the RHS \[4x=1.8-0.2\]\[4x=1.6\] Divide both sides by \(4\) \[\frac{4x}{4}=\frac{1.6}{4}\]\[x=0.4\]Do you understand now, @jennydoe?